# DualXDA: Towards Sparse, Efficient and Explainable Data Attribution in Large AI Models

### v2

@@ 3.1 DualDA: Using Support Vector Machines for Sparse and Efficient Data Attribution
This subsection presents a step-by-step derivation of DualDA attributions, with a graphical overview provided in Figure 1. DualDA uses the soft-margin multiclass SVM formulation introduced by Crammer and Singer (Crammer & Singer, 2001) as a surrogate model. The contribution of each training sample to the output of this surrogate model can be determined in a straightforward way, which we examine in detail below.
@@
Furthermore, we can align our surrogate-based approach with the framework of gradient-based approaches. Applying a generalized version of Influence Functions outlined in Corollary 1 of (Naujoks et al., 2024), we can investigate how the optimally derived in Equation 1 changes, under an infinitesimal downweighting of the contribution of datapoint and how this in turn affects the class score . This influence value aligns with the attribution calculated by DualDA, for datapoints that do not lie exactly on the margin. We report this result in Equation 8 and refer the reader to Appendix C for the proof.
@@ 3.2 XDA: Mapping Data Attribution to the Feature Space
While DualDA constitutes scalable and practically applicable solution for DA, we argue that DA alone is not sufficiently informative on its own: DA informs about which training datapoints are influential for a prediction, yet not why they are influential: A single datapoint may be relevant for various reasons, e.g., due to background patterns, parts of objects, or other semantic features. DA, however, does not provide any information to distinguish between these sources of information and their influence, resulting in potentially ambiguous explanations. To resolve this lack of clarity, we combine data attribution with feature attribution, leveraging the structure of our DualDA attributions with the introduction of XDA in order to obtain additional degrees of detail in our combined DualXDA explanations. This integrated perspective allows for deeper, more interpretable explanations of model behavior. While in principle, any FA approach could be levied to further explain the attributions of DualDA we choose to integrate the principles of Layer-wise Relevance Propagation (LRP) (Bach et al., 2015) to achieve XDA because like DualDA, it leverages the model’s activations on its hidden layers. LRP is a feature attribution technique based on the decomposition of a neural network’s prediction by backward propagating relevance scores layer-by-layer from the output to the input, following a relevance conservation principle that maintains and preserves the total amount of relevance throughout the decomposition process. As such, LRP is effectively striving towards the same goal as DualDA: the conservative attribution of model outputs to units of interpretation. However, the crucial difference between the two approaches lies in the orthogonal selection of the components targeted for interpretation, i.e., individual test features for LRP and training samples in their totality for DualDA. In the image domain, explanations derived from LRP are commonly represented as a heatmap over the corresponding input.
@@
The algorithm then backpropagates this quantity from layer of the DualDA surrogate model backwards through the feature extractor towards the input sample as well as the training sample in two separate backward passes. This procedure results in a pair of feature attributions and which together explain how – in terms of related features – the influential support vector in particular is affecting the model in the prediction of 333Note that while we can derive such explanations in all layers , we will focus on input-layer explanations in pixel space with in this paper. . As such, XDA shares conceptual similarities with BiLRP (Eberle et al., 2022) and CRP (Achtibat et al., 2023), but differs by explaining the feature kernel of a surrogate SVM (rather than a deep similarity model, as in BiLRP) and by splitting propagation pathways across support vectors instead of latent concepts (as in CRP). Overall, XDA leverages the conservative properties of DualDA and LRP to achieve data attribution as well as feature attribution, combining the benefits of the two orthogonal approaches. The conservative property of XDA further implies that the generated feature attribution heatmaps are also conservative whereby the sum of XDA test feature attribution maps recovers the LRP feature attribution map, i.e. LRP explanation of for class at layer is approximately equal to .
@@ 4 Evaluating Data Attribution
To test the validity of DualDA, we compare the method against other state-of-the-art attribution methods introduced in Section 2 and defined in detail in Section D.1. Details about implementations are given in Section 4.3. We analyze three different vision models of varying sizes, each trained on a different dataset: a simple Convolutional Neural Network with 3 convolutional and 3 fully connected layers with ReLU non-linearities trained on the MNIST dataset (LeCun et al., 2010), ResNet-18 (He et al., 2016) trained on CIFAR-10 (Krizhevsky & Hinton, 2009) and ResNet-50 (He et al., 2016) trained on the Animals with Attributes 2 (AwA2) dataset (Xian et al., 2018) (training details can be found in Section D.2). For each dataset and model, we calculate the attribution score of all training datapoints for the predicted class for 2,000 randomly selected test points, using the attribution methods. We evaluate the quality of the derived attributions on seven different metrics outlined in Section 4.1, two of which are novel contributions (for detailed explanations including formulas we refer the reader to Appendix E.). In our study, we rely on quanda (Bareeva et al., 2024), a toolkit designed for the systematic evaluation of DA methods. The results of our experiments are presented in Section 4.2. We further present the resource consumption of each attribution method in Section 4.3 and compare the sparsity of the corresponding explanations in Section 4.4. In Section 4.5, we analyze how close the surrogate model matches the original model for our DualDA method, as well as the Representer points method. All experiments have been executed on a single NVIDIA A100 Tensor Core-GPU.
@@ 4.1 Evaluation Metrics
Two simple sanity checks are proposed by Hanawa et al. (Hanawa et al., 2021): The Identical Class Test relies on the assumption that the most positively important training points for a class decision should be of the same class as the model prediction. We therefore measure for all test points the proportion of the highest ranking training data points according the DA methods outputs belonging to the same class as predicted for the given test sample. The Identical Subclass Test further assesses whether attributions can distinguish between different subpopulations within the same class. Following (Hanawa et al., 2021), we therefore artificially group dissimilar classes together into a superclass. Given a predictor trained on this new classification setting, we check whether the DA methods can still attribute influence to training samples from the same true subclass, rather than just the broader grouped class set, as the test sample.
Furthermore, two metrics are derived from common downstream tasks of DA: Mislabeling Detection (Koh & Liang, 2017; Yeh et al., 2018; Pruthi et al., 2020), which aims to discover incorrectly labeled training samples, and Shortcut Detection (Koh & Liang, 2017; Hammoudeh & Lowd, 2022), which tries to identify misleading associations in training data that allow models to achieve accuracy through dataset-specific biases rather than generalizable relationships.
Finally, retraining-based metrics test the counterfactual validity of training data attributions by retraining models on subsets of the training data and measuring whether the attributions align with observed changes in model behaviour. The Linear Datamodeling Score (LDS) (Park et al., 2023) tests whether attribution scores can predict changes in model outputs when retrained on random subsets of data. This is achieved by measuring the rank correlation between predicted attribution scores and actual model outputs. Coreset Selection evaluates whether attributions can identify the most valuable training samples by retraining models only on the highest-attributed data points and measuring how well performance is maintained. Adversarial Data Pruning tests the opposite of coreset selection by removing the most highly attributed training samples and measuring how much performance degrades. Unlike Coreset Selection, which evaluates whether attributions can identify a representative training subset, Adversarial Data Pruning assesses whether attributions correctly identify samples that contain unique information.
@@ 4.2 Results
From the experiments, we can also see how the choice of the sparsity hyperparameter affects the performance of DualDA on various metrics. This is particularly pronounced on Shortcut Detection, where a small sparsity is preferable, likely because it is preferable to have many training samples included in the support vectors to find all spuriously correlated samples. Conversely for Coreset Selection, higher sparsity yields markedly better results, plausibly because a representative subsample of the data is selected in the sparsification process. We extensively analyze the effect of the sparsity hyperparameter over a wide numerical range in Section F.3.
@@ 4.3 Resource Consumption
Computation time and memory requirements are the main concerns for scaling DA approaches to big models and datasets. Early approaches such as LiSSA (Koh & Liang, 2017) require time to attribute a new test point, where denotes the number of model parameters. This is indeed much slower than the explanation time of DualDA, which requires only time, as is the size of the penultimate layer of the network and therefore (for the ResNet-50 model, we have and ). Other approaches such as Arnoldi (Schioppa et al., 2022), EK-FAC (Grosse et al., 2023) and TRAK (Park et al., 2023) provide efficient approximations, which come at the cost of precomputing and and creating an information cache. In Figure 4, we report caching times, cache sizes and local explanations times (excluding caching times) for evaluations on 2,000 datapoints per dataset. The total runtime on a single GPU per method and dataset, which is the sum of caching and explanation time, is used as the -coordinate in each plot of Figure 2.
For TRAK, Representer Points and EK-FAC, we rely on official code releases of the original publications. For LiSSA, we use the implementation from a different publication (Bae et al., 2022) as it is written in the PyTorch (Paszke et al., 2019) framework. Employing the same framework universally guarantees a just comparison between all methods. We have implemented GradCos, GradDot and TracIn ourselves, using the random projection trick explained in (Pruthi et al., 2020). For Arnoldi, we relied on the captum (Kokhlikyan et al., 2020) implementation. To implement DualDA in Python, we have modified the multiclass SVM implementation from the Python package scikit-learn (Pedregosa et al., 2011) and included it in our repository1.
@@
Explanation Time To a substantial degree the slowest method at inference is LiSSA. Even though we have only included the final fully-connected layer in the Hessian and gradient computations, LiSSA takes 35 minutes for the explanation of a single test point of the AwA2 dataset, requiring almost 7 weeks for explaining the entire test set of 2,000 datapoints. This is the minimal runtime that is achieved by following the authors’ suggestion (Koh & Liang, 2017) that the total number of iterations dedicated to the estimation of the inverse-Hessian-vector-product of Equation D.1 be greater than the training dataset size. Arnoldi, TracIn and GradCos are moderately slow, requiring more than one second per test point on the AwA2 dataset. DualDA is the fastest method overall across all datasets, only requiring between 0.2ms and 4ms per test sample. Compared to the slowest method, LiSSA, DualDA is times faster on MNIST, times faster on CIFAR-10, and times faster on AwA2. Compared to the most efficient Influence Functions method, DualDA provides a speedup of , depending on the dataset. When generating explanations using a GPU, higher sparsity does not significantly decrease explanation time since the required matrix multiplication is already effectively parallelized and other parts of the code dominate the runtime. In our experiments, we observed, however, that when generating explanations on CPU, high sparsity significantly improves the computation time for explanations.
@@ 4.4 Sparsity of Attributions
We analyze the distribution of attribution scores over the training set for various DA methods, exemplified on the AwA2 dataset. Similar patterns are observable across other datasets in our study. Our results are displayed in Figure 5. Initially, we examine the overall ratio of cumulative positive attribution versus negative attribution, displayed on the left side of the figure. As expected, most explainers – with the exception of GradDot and GradCos – exhibit greater positive than negative cumulative attribution, which aligns with the fact that explanations target the model’s predicted class for each test instance. This naturally results in stronger positive attribution for proponents compared to negative attribution for opponents. The proportion of negative to positive cumulative attribution varies significantly across methods. The LiSSA implementation of Influence Functions shows the lowest ratio at only 53.7%, while TRAK demonstrates the highest at 98.7%. For DualDA, we observe that the sparsity hyperparameter directly affects this ratio – increasing values of (i.e., increasing sparsity) bring the positive cumulative relevance increasingly closer to the negative cumulative relevance.
On the right side of Figure 5, we analyze the growth of the cumulative sum of both positive and negative attribution as a function of the fraction of most influential datapoints included, to assess the sparsity characteristics of each method. We begin by incorporating training samples with the largest positive or largest negative attribution values respectively. Methods with sparse attribution patterns show steep initial increases in this curve, whereas non-sparse methods exhibit more gradual progression. Our findings reveal that the sparsity hyperparameter of DualDA effectively controls the sparsity of the resulting explanations. With , the method reaches full attribution saturation when selecting only 0.1% of the training samples descendingly ranked by attribution. Similarly, GradDot and EK-FAC show rapid initial increases, indicating concentrated attribution among few training samples. However, their growth rate quickly diminishes and they are eventually surpassed by Representer Points, which demonstrates slightly higher sparsity than DualDA configured with . LiSSA, TracIn, and GradCos exhibit particularly low sparsity, while the remaining methods display moderate sparsity characteristics. When comparing the development of the cumulative sum between positive and negative attribution, we observe that DualDA maintains consistent sparsity across both types. In contrast, Arnoldi and EK-FAC show increased sparsity for negative attribution, while Representer Points and GradDot demonstrate significantly reduced sparsity for negative attribution compared to positive attribution.
@@ 4.5 Faithfulness of Surrogates
Following the assumption that sparse attributions are preferable due to their lower complexity (Colin et al., 2022), it is only natural to decide to take a non-sparse attribution method and subsequently sparsify its results. As Representer Points is a method similar in spirit to DualDA, achieving promising results in attribution quality, we test the feasibility of this approach by post-hoc and artificially sparsifying the Representer Points surrogate model. We first sort the absolute coefficients for all classes and datapoints . Afterwards, we set all coefficients except for those with highest absolute value to zero. The number of non-zero coefficients for each class is set to the number of support vectors defining a trained DualDA surrogate for (an overview over the number of support vectors for each class and dataset can be found in Figure F.7 in Appendix F).
@@ 5 XDA Experiments
XDA provides unprecedented insight into the interaction between training and test samples by simultaneously highlighting relevant features in both input domains, thereby illuminating the similarities and dissimilarities that drive their relationship. This approach transforms previously opaque attribution values into interpretable explanations, offers a more nuanced understanding of model predictions, and enables feature-level ablation of training or test samples to analyze and modify training dynamics. To demonstrate XDA’s capabilities, we apply the approach to a small convolutional network trained on the MNIST dataset, and a VGG-16 network trained on AwA2 and the ImageNet dataset (Deng et al., 2009) respectively. Within the LRP framework, different formulas, or “rules”, exist to distribute the relevance from the output back to the input space. For our experiments, we use the recommended rule composite EpsilonPlusFlat, as implemented in the zennit toolkit (Anders et al., 2021), which applies the flat rule for any linear first layer, the z-plus rule for all other convolutional layers and the epsilon rule for all other fully connected layers. For VGG-16, we additionally apply the flat rule for the first 10 layers for better legibility of the corresponding relevance heatmaps. An explicit description of the corresponding formulas is given in (Montavon et al., 2019).

### v3

@@ 3.1 DualDA: Using Support Vector Machines for Sparse and Efficient Data Attribution
This subsection presents a step-by-step derivation of DualDA attributions, with a graphical overview provided in Figure 1. DualDA uses the soft-margin multiclass SVM formulation introduced by Crammer and Singer (svm:CS) as a surrogate model. The contribution of each training sample to the output of this surrogate model can be determined in a straightforward way, which we examine in detail below.
@@
Furthermore, we can align our surrogate-based approach with the framework of gradient-based approaches. Applying a generalized version of Influence Functions outlined in Corollary 1 of (naujoks2024pinnfluence), we can investigate how the optimally derived in Equation 1 changes, under an infinitesimal downweighting of the contribution of datapoint and how this in turn affects the class score . This influence value aligns with the attribution calculated by DualDA, for datapoints that do not lie exactly on the margin. We report this result in Equation 8 and refer the reader to Appendix C for the proof.
@@ 3.2 XDA: Mapping Data Attribution to the Feature Space
While DualDA constitutes scalable and practically applicable solution for DA, we argue that DA alone is not sufficiently informative on its own: DA informs about which training datapoints are influential for a prediction, yet not why they are influential: A single datapoint may be relevant for various reasons, e.g., due to background patterns, parts of objects, or other semantic features. DA, however, does not provide any information to distinguish between these sources of information and their influence, resulting in potentially ambiguous explanations. To resolve this lack of clarity, we combine data attribution with feature attribution, leveraging the structure of our DualDA attributions with the introduction of XDA in order to obtain additional degrees of detail in our combined DualXDA explanations. This integrated perspective allows for deeper, more interpretable explanations of model behavior. While in principle, any FA approach could be levied to further explain the attributions of DualDA we choose to integrate the principles of Layer-wise Relevance Propagation (LRP) (misc:xai:LRP) to achieve XDA because like DualDA, it leverages the model’s activations on its hidden layers. LRP is a feature attribution technique based on the decomposition of a neural network’s prediction by backward propagating relevance scores layer-by-layer from the output to the input, following a relevance conservation principle that maintains and preserves the total amount of relevance throughout the decomposition process. As such, LRP is effectively striving towards the same goal as DualDA: the conservative attribution of model outputs to units of interpretation. However, the crucial difference between the two approaches lies in the orthogonal selection of the components targeted for interpretation, i.e., individual test features for LRP and training samples in their totality for DualDA. In the image domain, explanations derived from LRP are commonly represented as a heatmap over the corresponding input.
@@
The algorithm then backpropagates this quantity from layer of the DualDA surrogate model backwards through the feature extractor towards the input sample as well as the training sample in two separate backward passes. This procedure results in a pair of feature attributions and which together explain how – in terms of related features – the influential support vector in particular is affecting the model in the prediction of 333Note that while we can derive such explanations in all layers , we will focus on input-layer explanations in pixel space with in this paper. . As such, XDA shares conceptual similarities with BiLRP (eberle2022bilrp) and CRP (achtibat2023attribution), but differs by explaining the feature kernel of a surrogate SVM (rather than a deep similarity model, as in BiLRP) and by splitting propagation pathways across support vectors instead of latent concepts (as in CRP). Overall, XDA leverages the conservative properties of DualDA and LRP to achieve data attribution as well as feature attribution, combining the benefits of the two orthogonal approaches. The conservative property of XDA further implies that the generated feature attribution heatmaps are also conservative whereby the sum of XDA test feature attribution maps recovers the LRP feature attribution map, i.e. LRP explanation of for class at layer is approximately equal to .
@@ 4 Evaluating Data Attribution
To test the validity of DualDA, we compare the method against other state-of-the-art attribution methods introduced in Section 2 and defined in detail in Section D.1. Details about implementations are given in Section 4.3. We analyze three different vision models of varying sizes, each trained on a different dataset: a simple Convolutional Neural Network with 3 convolutional and 3 fully connected layers with ReLU non-linearities trained on the MNIST dataset (data:mnist), ResNet-18 (arch:resnet-he) trained on CIFAR-10 (data:cifar10) and ResNet-50 (arch:resnet-he) trained on the Animals with Attributes 2 (AwA2) dataset (xian2018zero) (training details can be found in Section D.2). For each dataset and model, we calculate the attribution score of all training datapoints for the predicted class for 2,000 randomly selected test points, using the attribution methods. We evaluate the quality of the derived attributions on seven different metrics outlined in Section 4.1, two of which are novel contributions. For detailed explanations of evaluation metrics, including formulas, we refer the reader to Appendix E. In our study, we rely on quanda (bareeva2024quandainterpretabilitytoolkittraining), a toolkit designed for the systematic evaluation of DA methods. The results of our experiments are presented in Section 4.2. We further present the resource consumption of each attribution method in Section 4.3 and compare the sparsity of the corresponding explanations in Section 4.4. In Section 4.5, we analyze how close the surrogate model matches the original model for our DualDA method, as well as the Representer points method. All experiments have been executed on a single NVIDIA A100 Tensor Core-GPU.
@@ 4.1 Evaluation Metrics
Two simple sanity checks are proposed by Hanawa et al. (evaluation:hanawa): The Identical Class Test relies on the assumption that the most positively important training points for a class decision should be of the same class as the model prediction. We therefore measure for all test points the proportion of the highest ranking training data points according the DA methods outputs belonging to the same class as predicted for the given test sample. The Identical Subclass Test further assesses whether attributions can distinguish between different subpopulations within the same class. Following (evaluation:hanawa), we therefore artificially group dissimilar classes together into a superclass. Given a predictor trained on this new classification setting, we check whether the DA methods can still attribute influence to training samples from the same true subclass, rather than just the broader grouped class set, as the test sample.
Furthermore, two metrics are derived from common downstream tasks of DA: Mislabeling Detection (method:koh_influence; method:representer; method:trackin), which aims to discover incorrectly labeled training samples, and Shortcut Detection (method:koh_influence; hammoudeh2022identifying), which tries to identify misleading associations in training data that allow models to achieve accuracy through dataset-specific biases rather than generalizable relationships.
Finally, retraining-based metrics test the counterfactual validity of training data attributions by retraining models on subsets of the training data and measuring whether the attributions align with observed changes in model behaviour. The Linear Datamodeling Score (LDS) (misc:TRAK) tests whether attribution scores can predict changes in model outputs when retrained on random subsets of data. This is achieved by measuring the rank correlation between predicted attribution scores and actual model outputs. Coreset Selection evaluates whether attributions can identify the most valuable training samples by retraining models only on the highest-attributed data points and measuring how well performance is maintained. Adversarial Data Pruning tests the opposite of coreset selection by removing the most highly attributed training samples and measuring how much performance degrades. Unlike Coreset Selection, which evaluates whether attributions can identify a representative training subset, Adversarial Data Pruning assesses whether attributions correctly identify samples that contain unique information.
@@ 4.2 Results
From the experiments, we can also see how the choice of the sparsity hyperparameter affects the performance of DualDA on various metrics. This is particularly pronounced on Shortcut Detection, where a small sparsity is preferable, likely because it is preferable to have many training samples included in the support vectors to find all spuriously correlated samples. Conversely for Coreset Selection, higher sparsity yields markedly better results, plausibly because a representative subsample of the data is selected in the sparsification process. We extensively analyze the effect of the sparsity hyperparameter over a wide numerical range in Section F.3. Based on these results, we recommend a default setting of for image classification tasks with convolutional neural networks, which corresponds to the primary experimental configuration considered in this work.
@@ 4.3 Resource Consumption
Computation time and memory requirements are the main concerns for scaling DA approaches to big models and datasets. Early approaches such as LiSSA (method:koh_influence) require time to attribute a new test point, where denotes the number of model parameters. This is indeed much slower than the explanation time of DualDA, which requires only time, as is the size of the penultimate layer of the network and therefore (for the ResNet-50 model, we have and ). Other approaches such as Arnoldi (schioppa2022scaling), EK-FAC (grosse2023studying) and TRAK (misc:TRAK) provide efficient approximations, which come at the cost of precomputing and and creating an information cache. In Figure 4, we report caching times, cache sizes and local explanations times (excluding caching times) for evaluations on 2,000 datapoints per dataset. The total runtime on a single GPU per method and dataset, which is the sum of caching and explanation time, is used as the -coordinate in each plot of Figure 2.
For TRAK, Representer Points and EK-FAC, we rely on official code releases of the original publications. For LiSSA, we use the implementation from a different publication (bae2022if) as it is written in the PyTorch (misc:pytorch) framework. Employing the same framework universally guarantees a just comparison between all methods. We have implemented GradCos, GradDot and TracIn ourselves, using the random projection trick explained in (method:trackin). For Arnoldi, we relied on the captum (kokhlikyan_captum_2020) implementation. To implement DualDA in Python, we have modified444Code will be made available upon acceptance. the multiclass SVM implementation from the Python package scikit-learn (scikit-learn).
@@
Explanation Time To a substantial degree the slowest method at inference is LiSSA. Even though we have only included the final fully-connected layer in the Hessian and gradient computations, LiSSA takes 35 minutes for the explanation of a single test point of the AwA2 dataset, requiring almost 7 weeks for explaining the entire test set of 2,000 datapoints. This is the minimal runtime that is achieved by following the authors’ suggestion (method:koh_influence) that the total number of iterations dedicated to the estimation of the inverse-Hessian-vector-product of Equation D.1 be greater than the training dataset size. Arnoldi, TracIn and GradCos are moderately slow, requiring more than one second per test point on the AwA2 dataset. DualDA is the fastest method overall across all datasets, only requiring between 0.2ms and 4ms per test sample. Compared to the slowest method, LiSSA, DualDA is times faster on MNIST, times faster on CIFAR-10, and times faster on AwA2. Compared to the most efficient Influence Functions method, DualDA provides a speedup of , depending on the dataset. When generating explanations using a GPU, higher sparsity does not significantly decrease explanation time since the required matrix multiplication is already effectively parallelized and other parts of the code dominate the runtime. In our experiments, we observed, however, that when generating explanations on CPU, high sparsity significantly improves the computation time for explanations.
@@ 4.4 Sparsity of Attributions
We analyze the distribution of attribution scores over the training set for various DA methods, exemplified on the AwA2 dataset. Comparable patterns emerge across all datasets examined in this study. From a practical standpoint, sparsity can be quantified by determining the number of training points required to account for a given percentage of total attribution for a specific test sample. In Figure 5, we evaluate the sparsity characteristics of each method by analyzing how the cumulative sum of absolute attributions grows as a function of the fraction of most influential datapoints included. Training samples are incorporated in descending order of absolute attribution value. Methods exhibiting sparse attribution patterns demonstrate steep initial increases in this cumulative curve, while non-sparse methods show more gradual growth. Our findings reveal that the sparsity hyperparameter of DualDA effectively controls the sparsity of the resulting explanations. With , the method reaches full attribution saturation when selecting only 0.2% of the training samples descendingly ranked by absolute attribution. EK-FAC exhibits faster accumulation than Arnoldi, TRAK, GradDot, and Representer Points. Although these latter methods initially surpass DualDA configured with low sparsity , their accumulation rate decreases more rapidly, enabling DualDA with low sparsity to still reach 90% cumulative absolute attribution quicker. TracIn, GradCos, and LiSSA demonstrate minimal sparsity, distributing similar absolute attribution values across numerous training samples.
@@ 4.5 Faithfulness of Surrogates
Following the assumption that sparse attributions are preferable due to their lower complexity (colin2022cannot), it is only natural to decide to take a non-sparse attribution method and subsequently sparsify its results. As Representer Points is a method similar in spirit to DualDA, achieving promising results in attribution quality, we test the feasibility of this approach by post-hoc and artificially sparsifying the Representer Points surrogate model. We first sort the absolute coefficients for all classes and datapoints . Afterwards, we set all coefficients except for those with highest absolute value to zero. The number of non-zero coefficients for each class is set to the number of support vectors defining a trained DualDA surrogate for (an overview over the number of support vectors for each class and dataset can be found in Figure F.7 in Appendix F).
@@ 5 XDA Experiments
XDA provides unprecedented insight into the interaction between training and test samples by simultaneously highlighting relevant features in both input domains, thereby illuminating the similarities and dissimilarities that drive their relationship. This approach transforms previously opaque attribution values into interpretable explanations, offers a more nuanced understanding of model predictions, and enables feature-level ablation of training or test samples to analyze and modify training dynamics. To demonstrate XDA’s capabilities, we apply the approach to a small convolutional network trained on the MNIST dataset, and a VGG-16 network trained on AwA2 and the ImageNet dataset (deng2009imagenet) respectively. Within the LRP framework, different formulas, or “rules”, exist to distribute the relevance from the output back to the input space. For our experiments, we use the recommended rule composite EpsilonPlusFlat, as implemented in the zennit toolkit (anders2021software), which applies the flat rule for any linear first layer, the z-plus rule for all other convolutional layers and the epsilon rule for all other fully connected layers. For VGG-16, we additionally apply the flat rule for the first 10 layers for better legibility of the corresponding relevance heatmaps. An explicit description of the corresponding formulas is given in (montavon2019layer).